Secret:

Reduce the fraction to lowest terms. 38 x^2yz^2/-19 xy^2z^3

Mercury:

take it one variable at a time and apply the exponent rule in general: (a^m)/(a^n) = a^(m-n) can you try to apply this logic to (x^2)/x = ?

Mercury:

remember that, if no exponent is stated, the exponent is assumed to be 1

Mercury:

applying the logic from earlier, subtract the exponents so x^2 / x = x since 2 - 1 = 1

Mercury:

can you try to use this same logic to simplify y / y^2 = ?

Secret:

So, y / y^2 = y?

Mercury:

1 - 2 = -1 not 1 so y/y^2 = y^(-1)

Mercury:

order is important

Secret:

Oh

Mercury:

z^2/z^3 = ? remember, it's top exponent minus bottom exponent

Secret:

So that will be z?

Secret:

Wait no -z?

Mercury:

order is important 2 - 3 is -1 so z^(-1)

Secret:

Oh

Mercury:

now all that's left is 38/-19 which is just -2 putting everything together: 2xy^(-1) * z^(-1) if you have variables negative exponents, move them to the denominator therefore -2x/(y*z) = your solution

Secret:

Oh, ok! Thx